OpenStudy (anonymous):

what is the simplified form of (2x+4/x-2) - (x+3/x-2)

OpenStudy (cwrw238):

2x + 4 - (x + 3) ------------- x-2 simplify the top bit

OpenStudy (anonymous):

so would it be 3x-1/ x-2?

OpenStudy (anonymous):

i meant 3x+7/x-2

OpenStudy (cwrw238):

nope 2x - x = ? 4 - 3 = ?

OpenStudy (cwrw238):

the top simplifies to 2x + 4 - x - 3 = x + 1

OpenStudy (anonymous):

so it would be x+1/x-2

OpenStudy (cwrw238):

right